OK I have been stumped on this and skipped it but can you do it and see my answer matches because there is not a selection for mine. HERE IT IS---- (An algebra teacher drove by a farmyard full of chickens and pigs. The teacher happened to notice that there were a total of 100 heads and 270 legs. How many chickens were there? How many pigs were there?)
Answer by Alan3354(69443)

algebra teacher drove by a farmyard full of chickens and pigs. The teacher happened to notice that there were a total of 100 heads and 270 legs. How many chickens were there? How many pigs were there?)
------------
c + p = 100
2c + 4p = 270 --> c + 2p = 135
-------------
c + p = 100
c + 2p = 135
---------------- Subtract
-p = -35
--> 35 pigs
65 chickens